Vista Vacuum Company has the following production information for the month of March. All materials are added at the beginning of the manufacturing process.   

Units

  • Beginning inventory of 4,400 units that are 100 percent complete for materials and 28 percent complete for conversion.
  • 15,100 units started during the period.
  • Ending inventory of 3,800 units that are 14 percent complete for conversion.

Manufacturing Costs

  • Beginning inventory was $19,500 ($9,100 materials and $10,400 conversion costs).
  • Costs added during the month were $30,000 for materials and $54,600 for conversion ($27,600 labor and $27,000 applied overhead).

Assume the company uses Weighted-Average Method.

Required:
1.
Calculate the number of equivalent units of production for materials and conversion for March.
2. Calculate the cost per equivalent unit for materials and conversion for March.
3. Determine the costs to be assigned to the units transferred out and the units still in process.
Answer #1

Answers

  • Requirement 1

Physical units

Material

Conversion

Units completed

15700

15700

15700

Ending Inventory

3800

3800

532

TOTAL

19500

19,500

16,232

--Working

EUP - Weighted Average Method

Units

% Material

EUP Materials

% Conversion

EUP - Conversion

Units TRANSFERRED

15,700

100%

15,700

100%

15,700

Units of ENDING WIP

3,800

100%

3,800

14%

532

Equivalent Units of Production

19,500

16,232

  • Requirement 2

Material

Conversion

Cost per Equivalent units

$              2.00513

$                                      4.00444

--Working:

COST per EUP

Material

Conversion

Cost of Beginning WIP

$             9,100.00

$                                  10,400.00

Cost incurred during the period

$          30,000.00

$                                  54,600.00

Total Costs

Costs

$          39,100.00

Costs

$                                  65,000.00

Equivalent units of production

EUP

19,500

EUP

16,232

Cost per EUP

$              2.00513

$                                      4.00444

  • Requirement 3

Direct material

Conversion cost

Total

Units Completed

$                   31,481

$          62,870

$                  94,350

Ending Inventory

$                      7,619

$            2,130

$                    9,750

Total Cost accounted for

$                   39,100

$          65,000

$               104,100

--Working

Direct material

Conversion cost

Units Completed

=15700*2.00513

=15700*4.00444

Ending Inventory

=3800*2.00513

=532*4.00444